Operations Report – February 13th



Crew 275 Operations Report 13Feb2023
SOL: 1

Name of person filing report: Quentin ROYER
Non-nominal systems:
• Toilets
• Upper Deck outside weather station
• Spacesuit n°6

Notes on non-nominal systems:
• Weather station: the screen is on, but does not display any value and says “configure wifi for more detailed forecast”, and there is a red cross on the WiFi logo at the bottom left of the screen. How can we fix this?
• Spacesuit n°6: the collar is linked to the backpack with the help of only one fixation, instead of two on the other suits. Therefore this suit was not used today. There is a spare collar in the shelf in the EVA preparation room, do we have the permission to repair the spacesuit with it?

Spirit rover used: Yes
• Hours: 208.9
• Beginning charge: 100%
• Ending charge: 91%
• Currently charging: Yes

Opportunity rover used: Yes
• Hours: 113.0
• Beginning charge: 100%
• Ending charge: 92%
• Currently charging: Yes

Curiosity rover used: Yes
• Hours: 220.8
• Beginning charge: 100%
• Ending charge: 92%
• Currently charging: Yes

Perseverance rover used: Yes
• Hours: 255.4
• Beginning charge: 100%
• Ending charge: 88%
• Currently charging: Yes

General notes on rovers: Spirit and Opportunity were used during the morning EVA. Perseverance and Curiosity were used during the afternoon EVA.

Summary of Hab operations: The crew finished settling in the Hab, conducted the two training EVAs and started to set up the first experiments. An inventory of food, hygiene and medication items was performed All operations were nominal.
• WATER USE: 26.24 gallons
• Water: 458.41 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: No internet was used during the day before the opening of COMMS. We are using Hughes Net (Outpost 2g and Outpost 5g), and Starlink (MDRScampus-Guest).

Summary of suits and radios: Spacesuit n°6 is reported as non-nominal. The other suits are nominal. Radios were charged after the EA and are now all charged off-charger as the green-light was blinking. All radios were nominal.

Summary of GreenHab operations: The Crew botanist set up the aquarium for the aquaponics experiment. He then harvested tomato and basil.
• WATER USE: 21.78 gallons
• Heater: On
• Supplemental light: From 10 pm to 2 am
• Harvest: 223 g of tomato – 3 g of basil

Summary of Science Dome operations: An inventory was performed by the Crew Scientist.
• Dual split: Heat / From 10 pm to 7 am

Summary of RAM operations: An inventory was performed by the Crew Engineer.

Summary of any observatory issues: The Robotic Observatory was repaired and has been reported as operational by Mission Support. Operations are planned but the Crew astronomer for the next days, we will keep Mission support updated with regard to this.

Summary of health and safety issues: Small cut on the finger of a crew member.

Questions, concerns and requests to Mission Support:
• Is it possible to reduce / cut the fan that supplies air in the Commander / Executive Officer’s rooms?
• Could you remind us of the internet networks we are allowed to use, and their availability time slots?
• How can we monitor our internet consumption on Hughes Net and Starlink?
• What is the limit of internet consumption each day / during the mission?
• Where is the charger for the power drill that is located in the RAM?

Copyright © The Mars Society. All rights reserved. | Main Site